OpenStudy (anonymous):

Given that P(A) = 0.4, P(B) = 0.90, and P(A and B) = 0.30, determine P(A|B) 0.23 0.75 0.36 0.33

OpenStudy (mathmate):

P(A|B)=P(A intersect B) / P(B) So you have all the numbers needed to calculate P(A|B).

OpenStudy (anonymous):

still a little confused

OpenStudy (anonymous):

\[P(A|B)=\frac{P(A\cap B)}{P(B)}\]

OpenStudy (anonymous):

in your case \(P(A\cap B)=0.3,P(B)=0.9\)

OpenStudy (anonymous):

thats not an option up there

OpenStudy (jim766):

P(A|B) = P(A and B)/P(B) = 0.3/0.9 = 1/3

OpenStudy (jim766):

1/3 = .33333333333